# “Quantum Geometric Nesting” and Solvable Model Flat-Band Systems

### v1

@@ III Solvable ground state
Note that by doing a global gauge transformation $\hat{\gamma}\rightarrow\hat{\gamma}\mathrm{e}^{-\mathrm{i}\theta/2}$, we are able to transform ^$\hat{\mathcal{E}}(\theta)\rightarrow\hat{\mathcal{E}}(\theta=0)$ and thus $|\Psi_{\text{SC}}(\theta)\rangle\rightarrow|\Psi_{\text{SC}}(\theta=0)\rangle$, while keeping $\hat{S}^{(I)}_{\bm{R}}$ unchanged. Thus, all $|\Psi_{\text{SC}}(\theta)\rangle$ with different $\theta$ are simultaneous eigenstate of $\hat{S}^{(I)}_{\bm{R}}$ with the same eigenvalue. The corresponding ground state of $\hat{H}_{\text{int}}$ with a fixed (even) particle number $\mathsf{N}_{e}$ can be constructed as
@@
In the p-h case, the ground state of $\hat{H}_{\text{int}}$ can be similarly obtained by occupying any integer number of pseudo-bands of $\hat{\mathcal{E}}$ that are separated from other bands by finite gaps, since this gives rise to a non-degenerate  Note (3) and generically non-zero eigenvalue $\Phi_{0}$ within the corresponding particle number sector. It should be noted that when $\bm{Q}\neq 0$, $\hat{\mathcal{E}}$ has a folded BZ since it breaks the translation symmetry of the original electronic structure. Diagonalizing this pseudo-Hamiltonian then gives us $\mathsf{M}\mathsf{N}_{\text{flat}}$ pseudo-bands, where $\mathsf{M}$ is the smallest positive integer that makes $\mathsf{M}\bm{Q}=0$ modulo the reciprocal lattice. In contrast to the p-p case, here $\theta$ only cycles the pseudo-energies, permuting the equivalent translation-symmetry-breaking groundstates.
@@ IV.1 Few-body excitations
This is easier said than done for typical interacting Hamiltonians, where the commutator of a one-particle operator, $\gamma_{a}^{\dagger}$, with the Hamiltonian produces higher order terms such as $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$, and the commutator of many-particle operators produces still higher order terms in a hierarchy that never closes. However, because $|\Psi\rangle$ is a zero-mode of all $\left(\hat{S}^{(I)}_{\bm{R}}-\langle\hat{S}^{(I)}_{\bm{R}}\rangle\right)$ within $\mathcal{H}_{\text{flat}}$, sectors with e.g. $\gamma_{a}^{\dagger}$ and $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$ excitations cannot mix Kang and Vafek (2019); Bernevig et al. (2021); Herzog-Arbeitman et al. (2022a); Vafek and Kang (2021) and thus the composition does not changed after the commutation. For example, for two-particle (Cooper pair) excitations,
@@
where $\bm{k}^{\pm}\equiv\bm{P}/2\pm\bm{k}$. The commutation relations for single-particle/hole, two-hole, and particle-hole (exciton) excitations can be similarly derived (see Sec. IV A-C of SM for computationally feasible expressions in terms of parameters in the Hamiltonian). By diagonalizing the scattering matrix $\Gamma$, one can compute the corresponding eigenstates and spectrum of an excitation Bernevig et al. (2021); Herzog-Arbeitman et al. (2022a). For example, the charge $+1$ excitation spectrum only involves diagonalizing an effective single-particle Hamiltonian. These excitations can be computed with polynomial scaling in the system size, rather than exponentially as in most interacting systems, and in this sense are solvable. For certain cases when the form of interaction is simple (e.g. attractive Hubbard Herzog-Arbeitman et al. (2022a)), the quasi-particle spectrum can be even analytically solvable.
@@ V.1 Spinful time-reversal symmetric flat bands
Then one can verify $N^{\text{p-p},\bm{Q}=\bm{0}}_{\mu s,\nu s^{\prime}}\equiv\sigma^{y}_{ss^{\prime}}\mathbbm{1}_{\mu\nu}$ satisfies the perfect QGN condition in the p-p channel with $\bm{Q}=\bm{0}$. Solvable Hamiltonians can thus be constructed with $B^{(I)}_{\mu s,\nu s^{\prime}}$ matrices in Eq. 15 taking the form $\sigma^{\mathsf{i}=0,x,y,z}_{ss^{\prime}}\tilde{B}^{(I)}_{\mu\nu}$, where $\sigma^{\mathsf{i}}$ are spin Pauli matrices and $\tilde{B}^{(I)}$ are arbitrary symmetric (for $\mathsf{i}=x,y,z$) or antisymmetric (for $\mathsf{i}=0$) hermitian matrices for the orbital indices. This includes a huge class of possible interactions, e.g. the attractive Hubbard interactions on each site and orbital, whose solvability was pointed out in Refs. Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022a) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al. (2023); Liang et al. (2017); Verma et al. (2021); Hazra et al. (2018); Xie et al. (2020); Herzog-Arbeitman et al. (2022b); Huhtinen et al. (2022); Herzog-Arbeitman et al. (2022a).
@@ V.2 Chiral symmetric flat bands
A less obvious example is associated with chiral symmetry. Specifically, when the lattice is bipartite and there are degenerate flat bands, one can verify that $N^{\text{p-h},\bm{Q}=\bm{0}}_{\mu s,\nu s^{\prime}}=\delta_{\mu\nu}(-1)^{\mu}\tilde{N}_{ss^{\prime}}$ is a nesting matrix for the p-h channel at $\bm{Q}=\bm{0}$, where the parity of the orbital index $\mu$ represents the sublattice group, and $\tilde{N}$ is an arbitrary hermitian matrix associated with the conserved flavor index $s$ (e.g. spin and/or valley). The condensation of the corresponding OP then leads to symmetry breaking of the conserved flavor symmetry with a pattern specified by $\tilde{N}$ (e.g. spin/valley polarization, inter-valley coherence, …). This is exactly the case extensively discussed in the studies on magic-angle twisted bilayer graphene in its chiral-flat limit Kang and Vafek (2019); Vafek and Kang (2021); Bultinck et al. (2020); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021). We also note that there is a large class of electronic structures that feature flat bands protected by chiral symmetry Călugăru et al. (2022) and are approximately realized in materials Regnault et al. (2022).
@@ V.3 Non-symmetry-related cases: An engineered model for pair-density-wave
with $\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\nu}=\sum_{\nu}\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\mu}P_{\mu\nu}(\bm{k})$ the projection of the orbital operators onto the flat bands (see Eq. (9)). Projecting $\hat{H}_{\text{int}}$ in Eq. (31) to the flat band Hilbert space results in trivial one-body (Hartree-Fock-type) terms that act as a chemical potential on the flat bands. The triviality of the one-body terms generalizes the “uniform pairing condition” Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022a) utilized in attractive Hubbard models (see Sec. II D of the SM).
Adopting the methods in Sec. IV, we can exactly solve the few-body excitations in an artificial setup by setting the $W\to 0$ and $\Delta\to\infty$ with the quantum geometry $U_{\mu n}(\bm{k})$ remaining fixed. First, we consider the charge ±$\pm 1$ fermionic quasi-particles above the ground states. We find that these excitations are spin-degenerate, gapped, and flat, with energy $E^{(\pm 1)}(\mathbf{p})=3V$. Next, we consider a Cooper pair excitation. This spectrum is also analytically solvable since the two-particle scattering matrix can be diagonalized by generalizing the approach of Ref. Herzog-Arbeitman et al. (2022a) to extract a low-rank Hamiltonian for the pair binding energies. We find that there are two bound modes - the Goldstone mode of the PDW condensate, and a higher-energy Leggett mode - below the flat and macroscopically degenerate two-particle continuum at $6V$. The Goldstone mode has dispersion
@@
which applies as long as $\Delta\sim M\gg V\gg W\sim t^{2}/M$. In the further limit $V\gg t$, we see that $T_{c}\gg W^{2}/V$, the conventional strong coupling scaling for single-band systems. This observation further corroborates that the phase stiffness originates from the quantum geometry of the band instead of the (single-particle) kinetic energies. This example (along with similar recent efforts in Refs. Jiang and Barlas (2023); Chen and Huang (2023)) is based on the geometric structure of the flat bands and is distinguished by its analytically controlled, microscopic approach. QGN thus points to a fresh direction in the pursuit of PDW phases, intrinsically distinct from the existing literature Han and Kivelson (2022); Wu et al. (2023a); Liu and Han (2023); Wu et al. (2023b); Jin et al. (2022); Han et al. (2020); Setty et al. (2023); Haldane et al. (2023).
@@ VI Possible Extensions
which realizes an obstructed atomic limit Song et al. (2020) with lower-bounded quantum geometry Herzog-Arbeitman et al. (2022b); Espinosa-Champo and Naumis (2023). As discussed in the previous section, the time-reversal symmetry and spin conservation imply perfect QGN in the p-p channel, suggesting a singlet $s$-wave superconducting groundstate Herzog-Arbeitman et al. (2022a). However, we find that this result can be promoted to a singlet $p$-wave OP
@@
by choosing the interaction taking the form $H_{\text{int}}=U\sum_{R}\hat{S}^{\dagger}_{\bm{R}}\hat{S}_{\bm{R}}\geq 0$ where $\hat{S}_{\bm{R}}$ again satisfies $[\hat{P}\hat{S}_{\bm{R}}\hat{P},\hat{O}^{\text{p-p},\dagger}]=0$ but is non-hermitian. This generalizes the form assumed in Eq. 12 (see Sec. VI of SM for the explicit construction which guarantees trivial one-body terms). Remarkably, zero-energy ground states can still be exactly constructed according to $(\hat{O}^{\text{p-p},\dagger})^{n}\ket{0}$ and proven to be unique using a corollary of Lieb’s well-known theorem Lieb (1989). Although the excitations in this construction are not exactly solvable, we can prove (see Sec. VI of SM) that the charge ±$\pm 1$ excitation spectrum is gapless, as expected for a $p$-wave ground state. This example demonstrates that the possible construction schemes of “ideal” interacting Hamiltonians for systems with QGN have not been exhausted, which we leave for future explorations.
Moreover, it remains to be investigated whether QGN is the necessary condition for a set of flat bands to admit solvable strongly coupled interacting Hamiltonians. Our work advances the exploration Salerno et al. (2023); Huang et al. (2022); Yu et al. (2023); Kashihara et al. (2023); Herzog-Arbeitman et al. (2022); Soldini et al. (2023); Jiang et al. (2023); Yu et al. (2023); Peralta Gavensky et al. (2023); Setty et al. (2023, 2023) of many-body phases enabled by topology and quantum geometry making use of the strongly coupled flat band limit, but it also important to explore how QGN can be combined with Fermi surface nesting for dispersive systems featuring significant quantum geometry.
Acknowledgement: We would like to thank Erez Berg, Daniel Parker, Junkai Dong, Ashvin Vishwanath, Chaitanya Murthy, Biao Lian, Ben Feldman, Oskar Vafek, Jiabin Yu, Zhi-da Song, and Päivi Törmä for helpful discussions. ZH was funded, in part, by a QuantEmX grant from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616. BAB was primarily supported by the the Simons Investigator Grant No. 404513. JHA was supported by a Hertz Fellowship. BAB and JHA received additional support from the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, Office of Naval Research (ONR Grant No. N00014-20-1-2303), BSF Israel US foundation No. 2018226 and NSF-MERSEC DMR-2011750, Princeton Global Scholar and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No 101017733 and from the European Research Council (ERC). SAK and ZH were supported, in part, by the Department of Energy, Office of Basic Energy Sciences, under contract DE-AC0276SF00515 at Stanford.

### v2

@@ III Solvable ground state
Note that by doing a global gauge transformation $\hat{\gamma}\rightarrow\hat{\gamma}\mathrm{e}^{-\mathrm{i}\theta/2}$, we are able to transform ^$\hat{\mathcal{E}}(\theta)\rightarrow\hat{\mathcal{E}}(\theta=0)$ and thus $|\Psi_{\text{SC}}(\theta)\rangle\rightarrow|\Psi_{\text{SC}}(\theta=0)\rangle$, while keeping $\hat{S}^{(I)}_{\bm{R}}$ unchanged. Thus, all $|\Psi_{\text{SC}}(\theta)\rangle$ with different $\theta$ are simultaneous eigenstate of $\hat{S}^{(I)}_{\bm{R}}$ with the same eigenvalue. The corresponding ground state of $\hat{H}_{\text{int}}$ with an arbitrary fixed (even) particle number $\mathsf{N}_{e}$ can be constructed as
@@
In the p-h case, the ground state of $\hat{H}_{\text{int}}$ can be similarly obtained by occupying any integer number of pseudo-bands of $\hat{\mathcal{E}}$ that are separated from other bands by finite gaps, since this gives rise to a non-degenerate  Note (3) and generically non-zero eigenvalue $\Phi_{0}$ within the corresponding particle number sector. This implies that the system is solvable only at certain rational filling fractions of the flat bands with specific values determined by the OP. It should be noted that when $\bm{Q}\neq 0$, $\hat{\mathcal{E}}$ has a folded BZ since it breaks the translation symmetry of the original electronic structure. Diagonalizing this pseudo-Hamiltonian then gives us $\mathsf{M}\mathsf{N}_{\text{flat}}$ pseudo-bands, where $\mathsf{M}$ is the smallest positive integer that makes $\mathsf{M}\bm{Q}=0$ modulo the reciprocal lattice. In contrast to the p-p case, here $\theta$ only cycles the pseudo-energies, permuting the equivalent translation-symmetry-breaking groundstates.
@@ IV.1 Few-body excitations
This is easier said than done for typical interacting Hamiltonians, where the commutator of a one-particle operator, $\gamma_{a}^{\dagger}$, with the Hamiltonian produces higher order terms such as $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$, and the commutator of many-particle operators produces still higher order terms in a hierarchy that never closes. However, because $|\Psi\rangle$ is a zero-mode of all $\left(\hat{S}^{(I)}_{\bm{R}}-\langle\hat{S}^{(I)}_{\bm{R}}\rangle\right)$ within $\mathcal{H}_{\text{flat}}$, sectors with e.g. $\gamma_{a}^{\dagger}$ and $\gamma_{b}^{\dagger}\gamma_{c}^{\dagger}\gamma_{d}$ excitations cannot mix Kang and Vafek (2019); Bernevig et al. (2021); Herzog-Arbeitman et al. (2022); Vafek and Kang (2021) and thus the composition does not changed after the commutation. For example, for two-particle (Cooper pair) excitations,
@@
where $\bm{k}^{\pm}\equiv\bm{P}/2\pm\bm{k}$. The commutation relations for single-particle/hole, two-hole, and particle-hole (exciton) excitations can be similarly derived (see Sec. IV A-C of SM for computationally feasible expressions in terms of parameters in the Hamiltonian). By diagonalizing the scattering matrix $\Gamma$, one can compute the corresponding eigenstates and spectrum of an excitation Bernevig et al. (2021); Herzog-Arbeitman et al. (2022). For example, the charge $+1$ excitation spectrum only involves diagonalizing an effective single-particle Hamiltonian. These excitations can be computed with polynomial scaling in the system size, rather than exponentially as in most interacting systems, and in this sense are solvable. For certain cases when the form of interaction is simple (e.g. attractive Hubbard Herzog-Arbeitman et al. (2022)), the quasi-particle spectrum can be even analytically solvable.
@@ V.1 Spinful time-reversal symmetric flat bands
Then one can verify $N^{\text{p-p},\bm{Q}=\bm{0}}_{\mu s,\nu s^{\prime}}\equiv\sigma^{y}_{ss^{\prime}}\mathbbm{1}_{\mu\nu}$ satisfies the perfect QGN condition in the p-p channel with $\bm{Q}=\bm{0}$. Solvable Hamiltonians can thus be constructed with $B^{(I)}_{\mu s,\nu s^{\prime}}$ matrices in Eq. 15 taking the form $\sigma^{\mathsf{i}=0,x,y,z}_{ss^{\prime}}\tilde{B}^{(I)}_{\mu\nu}$, where $\sigma^{\mathsf{i}}$ are spin Pauli matrices and $\tilde{B}^{(I)}$ are arbitrary symmetric (for $\mathsf{i}=x,y,z$) or antisymmetric (for $\mathsf{i}=0$) hermitian matrices for the orbital indices. This includes a huge class of possible interactions, e.g. the attractive Hubbard interactions on each site and orbital, whose solvability was pointed out in Refs. Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022) and where the superfluid stiffness is shown to fundamentally depend on the flat-band quantum geometry and topology in several previous works Hofmann et al. (2023); Liang et al. (2017a); Verma et al. (2021); Hazra et al. (2018); Xie et al. (2020); Herzog-Arbeitman et al. (2022); Huhtinen et al. (2022); Herzog-Arbeitman et al. (2022).
@@ V.2 Chiral symmetric flat bands
A less obvious example is associated with chiral symmetry. Specifically, when the lattice is bipartite and there are degenerate flat bands, one can verify that $N^{\text{p-h},\bm{Q}=\bm{0}}_{\mu s,\nu s^{\prime}}=\delta_{\mu\nu}(-1)^{\mu}\tilde{N}_{ss^{\prime}}$ is a nesting matrix for the p-h channel at $\bm{Q}=\bm{0}$, where the parity of the orbital index $\mu$ represents the sublattice group, and $\tilde{N}$ is an arbitrary hermitian matrix associated with the conserved flavor index $s$ (e.g. spin and/or valley). The condensation of the corresponding OP then leads to symmetry breaking of the conserved flavor symmetry with a pattern specified by $\tilde{N}$ (e.g. spin/valley polarization, inter-valley coherence, …). This is exactly the case extensively discussed in the studies on magic-angle twisted bilayer graphene in its chiral-flat limit Kang and Vafek (2019); Vafek and Kang (2021); Bultinck et al. (2020); Lian et al. (2021); Bernevig et al. (2021); Ledwith et al. (2021). In the language of QGN, it is now clear that these proposed solvable interacting Hamiltonians belong to a larger class, which is systematically constructible using the method in Sec. II for each possible ordering. We also note that there is a large class of electronic structures that feature flat bands protected by chiral symmetry Călugăru et al. (2022) and are approximately realized in materials Regnault et al. (2022).
@@ V.3 Non-symmetry-related cases: An engineered model for pair-density-wave
with $\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\nu}=\sum_{\nu}\hat{\bar{c}}^{\dagger}_{\bm{k}+\bm{Q},\mu}P_{\mu\nu}(\bm{k})$ the projection of the orbital operators onto the flat bands (see Eq. (9)). Projecting $\hat{H}_{\text{int}}$ in Eq. (31) to the flat band Hilbert space results in trivial one-body (Hartree-Fock-type) terms that act as a chemical potential on the flat bands. The triviality of the one-body terms generalizes the “uniform pairing condition” Tovmasyan et al. (2016); Herzog-Arbeitman et al. (2022) utilized in attractive Hubbard models (see Sec. II D of the SM).
Adopting the methods in Sec. IV, we can exactly solve the few-body excitations in an artificial setup by setting the $W\to 0$ and $\Delta\to\infty$ with the quantum geometry $U_{\mu n}(\bm{k})$ remaining fixed. First, we consider the charge ±$\pm 1$ fermionic quasi-particles above the ground states. We find that these excitations are spin-degenerate, gapped, and flat, with energy $E^{(\pm 1)}(\mathbf{p})=3V$. Next, we consider a Cooper pair excitation. This spectrum is also analytically solvable since the two-particle scattering matrix can be diagonalized by generalizing the approach of Ref. Herzog-Arbeitman et al. (2022) to extract a low-rank Hamiltonian for the pair binding energies. We find that there are two bound modes - the Goldstone mode of the PDW condensate, and a higher-energy Leggett mode - below the flat and macroscopically degenerate two-particle continuum at $6V$. The Goldstone mode has dispersion
@@
which applies as long as $\Delta\sim M\gg V\gg W\sim t^{2}/M$. In the further limit $V\gg t$, we see that $T_{c}\gg W^{2}/V$, the conventional strong coupling scaling for single-band systems. This observation further corroborates that the phase stiffness originates from the quantum geometry of the band instead of the (single-particle) kinetic energies. This example (along with similar efforts in Refs. Liang et al. (2017b); Jiang and Barlas (2023); Chen and Huang (2023)) is based on the geometric structure of the flat bands and is distinguished by its analytically controlled, microscopic approach. QGN thus points to a fresh direction in the pursuit of PDW phases, intrinsically distinct from the existing literature Han and Kivelson (2022); Wu et al. (2023a); Liu and Han (2024); Wu et al. (2023b); Jin et al. (2022); Han et al. (2020); Setty et al. (2023b); Haldane et al. (2023).
@@ VI Possible Extensions
which realizes an obstructed atomic limit Song et al. (2020) with lower-bounded quantum geometry Herzog-Arbeitman et al. (2022); Espinosa-Champo and Naumis (2023). As discussed in the previous section, the time-reversal symmetry and spin conservation imply perfect QGN in the p-p channel, suggesting a singlet $s$-wave superconducting groundstate Herzog-Arbeitman et al. (2022). However, we find that this result can be promoted to a singlet $p$-wave OP
@@
by choosing the interaction taking the form $H_{\text{int}}=U\sum_{R}\hat{S}^{\dagger}_{R}\hat{S}_{R}\geq 0$ where $\hat{S}_{R}$ again satisfies $[\hat{P}\hat{S}_{R}\hat{P},\hat{O}^{\text{p-p},\dagger}]=0$ but is non-hermitian. Explicitly, we choose $\hat{S}_{R}$ to take the form
$\hat{S}_{R}$ $=\sum_{\alpha\beta,\sigma}\sum_{\Delta R=-2}^{2}s^{z}_{\sigma}\tau_{\alpha\beta}(\Delta R)\hat{c}^{\dagger}_{R+\Delta R,\alpha,\sigma}\hat{c}_{R,\beta,\sigma}$ (38)
where $s^{z}_{\uparrow/\downarrow}=\pm 1$ and $\tau(0)=\tau_{2}/4,\tau(\pm 1)=\mp i\tau_{0}/4,\tau(\pm 2)=(-\tau_{2}\pm i\tau_{3})/8$. This generalizes the form assumed in Eq. 12 (see Sec. VI of SM for the explicit construction which guarantees trivial one-body terms). Remarkably, zero-energy ground states can still be exactly constructed according to $(\hat{O}^{\text{p-p},\dagger})^{n}\ket{0}$ and proven to be unique using a corollary of Lieb’s well-known theorem Lieb (1989). Although the excitations in this construction are not exactly solvable, we can prove (see Sec. VI of SM) that the charge ±$\pm 1$ excitation spectrum is gapless, as expected for a $p$-wave ground state. This example demonstrates that the possible construction schemes of “ideal” interacting Hamiltonians for systems with QGN have not been exhausted, which we leave for future explorations.
Moreover, it remains to be investigated whether QGN is the necessary condition for a set of flat bands to admit solvable strongly coupled interacting Hamiltonians. It is interesting to explore whether new notions in quantum geometry Bouhon et al. (2023) can help discover more general conditions of QGN. Our work advances the exploration of many-body phases enabled by topology and quantum geometry making use of the strongly coupled flat band limit, but it also important to explore how QGN can be combined with Fermi surface nesting for dispersive systems featuring significant quantum geometry.
Acknowledgement: We would like to thank Erez Berg, Daniel Parker, Junkai Dong, Ashvin Vishwanath, Chaitanya Murthy, Biao Lian, Ben Feldman, Oskar Vafek, Jiabin Yu, Zhi-da Song, and Päivi Törmä for helpful discussions. We also thank anonymous reviewers for inspiring questions and comments. ZH was funded, in part, by a QuantEmX grant from ICAM and the Gordon and Betty Moore Foundation through Grant GBMF9616. BAB was primarily supported by the the Simons Investigator Grant No. 404513. JHA was supported by a Hertz Fellowship. BAB and JHA received additional support from the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, Office of Naval Research (ONR Grant No. N00014-20-1-2303), BSF Israel US foundation No. 2018226 and NSF-MERSEC DMR-2011750, Princeton Global Scholar and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No 101017733 and from the European Research Council (ERC). SAK and ZH were supported, in part, by the Department of Energy, Office of Basic Energy Sciences, under contract DE-AC0276SF00515 at Stanford.